### v1

# Steering Large Language Models for Machine Translation Personalization

Abstract
High-quality machine translation systems based on large language models (LLMs) have simplified the production of personalized translations reflecting specific stylistic constraints. However, these systems still struggle in settings where stylistic requirements are less explicit and might be harder to convey via prompting. We explore various strategies for personalizing LLM-generated translations in low-resource settings, focusing on the challenging literary translation domain. We explore prompting strategies and inference-time interventions for steering model generations towards a personalized style, and propose a contrastive framework exploiting latent concepts extracted from sparse autoencoders to identify salient personalization properties. Our results show that steering achieves strong personalization while preserving translation quality. We further examine the impact of steering on LLM representations, finding model layers with a relevant impact for personalization are impacted similarly by multi-shot prompting and our steering method, suggesting similar mechanism at play. [^1] Code available at DanielSc4/steering-for-personalization.
{CJK*}
UTF8gbsn
1 Introduction
Figure 1: We compare prompt-based approaches with steering techniques intervening on model internals for personalizing MT outputs in literary machine translation, employing MT quality metrics and style classifiers to disentangle the effect of steering on outputs fluency and personalization adequacy.
When we read a translated book, we do not simply read the story in a new language; we also experience the translator’s personal voice by means of their stylistic choices. Past efforts in the automatic translation of literary works have historically been constrained by the limited capabilities and flexibility of machine translation (MT) systems. The recent popularization of MT systems based on large language models (LLMs) has greatly improved their capacity in handling the long contexts typical of literary translations, but mimicking the creative and rich language that characterize the translators’ own style remains an open issue. In this context, several works explored the usage of prompting and tuning-based strategies to ensure translations are stylistically appropriate (Michel and Neubig, 2018; Wang et al., 2021). However, their influence on model internal representations is rarely explored, so that their impact is less controllable and often unpredictable. In this work, we adopt prompting methods as baselines and compare their effectiveness to steering methods proposed in interpretability literature. These techniques can be used to surgically intervene on LLMs’ intermediate representation to generate personalized translations when few examples are available. For this purpose, we employ the par3 dataset (Thai et al., 2022), which includes multiple human translations for novels translated to English from 7 typologically diverse languages.
We begin with preliminary assessments by verifying whether translators’ styles are discernible by automatic systems, finding that trained classifiers can distinguish writing styles with high accuracy, while the task is notoriously challenging for human annotators (Youyou et al., 2015; Flekova et al., 2016). We also find a simple prompting setting with in-context personalization examples to improve the style accuracy of LLM translation, suggesting personalized translation styles are reproducible. We connect the conditioning induced by prompting to the inner workings of the model, identifying activations with high discriminative capacity for style differences in intermediate model layers. We then propose a contrastive steering approach based on sparse autoencoders (SAEs, Cunningham et al., 2023) to condition model generations by upweighting sparse, interpretable latents at inference time. We validate the effectiveness of our method across three LLMs of various sizes on par3 novels, comparing our results with established prompting and steering methods.
Our results show that contrastive SAE steering is a promising approach for MT personalization, leading to translations that are not only more in line with general human translation features but also more aligned with the desired personalized style compared to other methods. Importantly, these results are achieved with no translation quality degradation according to established MT quality metrics. We conclude by comparing the impact of our method on model representations with the outcome of multi-shot prompting, finding that probes trained on prompt-conditioned activations can predict the effectiveness of SAE steering with high precision. These results confirm that tested prompting and steering techniques converge to similar solutions for conditioning model behavior, enabling future investigations into the mechanistic impact of prompting through the study of learned SAE latents and other interpretable components.
2 Related Work
Machine Translation of Literary Texts
The literary domain has historically been challenging for automatic MT systems due to their limited ability in handling rich linguistic and cultural contexts (Matusov, 2019) and their propensity to produce overly literal outputs (Guerberof-Arenas and Toral, 2022). Automatic literary translation has a long history dating back to pre-neural MT approaches (Voigt and Jurafsky, 2012; Toral and Way, 2015, 2018; Moorkens et al., 2018) with two recent dedicated evaluation campaigns (Wang et al., 2023b, 2024a). The advent of LLMs brought new opportunities in the processing of longer context for document-level translation (Wang et al., 2023a; Briakou et al., 2024; Wu et al., 2025a), but critical errors requiring human translator’s intervention nonetheless persist (Karpinska and Iyyer, 2023). In this work, we use the par3 dataset (Thai et al., 2022) containing multiple human translations of novels to evaluate MT personalization in the literary domain.
Personalization for Machine Translation
Advances in MT quality recently led to a growing interest in personalization approaches to ensure a consistent format and appropriate stylistic choices in model generations (Rabinovich et al., 2017; Lin et al., 2021). Previous approaches for controlling attributes such as formality (Sennrich et al., 2016; Niu et al., 2017; Nadejde et al., 2022) or gender (Vanmassenhove et al., 2018; Saunders and Byrne, 2020) typically required tuning existing models on pre-defined properties of interest, with few works attempting a real data-driven adaptation from unlabeled demonstrations (Michel and Neubig, 2018; Wang et al., 2021; Zhang et al., 2022). More recently, several studies employed prompting (Garcia and Firat, 2022; Sarti et al., 2023) or preference optimization from post-editing behavior (Lee et al., 2023; Berger et al., 2024) to render MT personalization more effective and data-efficient. In this work, we complement prompt results with steering approaches to personalize MT outputs using few user-provided examples.
Steering Language Model Generations
Steering approaches exploit the linear structure of LM activations (Mikolov et al., 2013; Chanin et al., 2024) to craft inference-time interventions for influencing model generations. These methods commonly employ contrastive sets of in-context demonstrations (Rimsky et al., 2024; Scalena et al., 2024) to map input properties to components such as vectors (Turner et al., 2024; Li et al., 2023), linear probes (Zou et al., 2025) or learned projections (Wu et al., 2024, 2025b). Sparse Autoencoders (SAEs) are another family of promising approaches for enabling fine-grained interventions in language models (Yun et al., 2021; Cunningham et al., 2023; Templeton et al., 2024). They are trained to decompose activations into approximately monosemantic features, offering a potentially interpretable basis for modifying model behavior. While interpreting their learned latents remains non-trivial (Marks et al., 2025), SAEs have proven effective for applying targeted interventions along specific linear directions (Chalnev et al., 2024; Zhao et al., 2025; Ferrando et al., 2025). However, most research on SAEs have so far focused on synthetic tasks or standard benchmarks, leaving their potential in real-world settings relatively underexplored.
3 Preliminaries
zh $\rightarrow$ en   道人道:「既如此,便隨你去來。」
H1 "Such being the case," the Taoist acquiesced, "I am ready to follow you, whenever you please to go."
H2 "Very good, I will go with you then," said the Taoist.
Gemma 2 2B "If that’s the case, then go ahead" said the Taoist.
Gemma 2 9B "If so, then I will go with you" the Taoist said
Llama 3.1 8B The Taoist said: "If you insist on going, then go ahead."
it $\rightarrow$ en   Sarà l’effetto dell’acqua del mare. Il mare ne fa di questi scherzi
H1 "The salt water must have done it. The sea plays funny tricks."
H2 "It must have been the effect of sea-water. The sea makes extraordinary changes."
Gemma 2 2B "It will be the effect of the sea water. The sea makes of these jokes."
Gemma 2 9B It will be the effect of the sea water. The sea plays these tricks.
Llama 3.1 8B It will be the effect of the sea water. The sea does things like this.
Table 1: zh $\rightarrow$ en and it $\rightarrow$ en examples for par3 segments translated by humans (H1, H2) and LLMs with zero-shot prompting. More examples in Appendix D.
Before testing the effectiveness of personalization strategies, we validate some key assumptions: i) Whether the personalized translation style is discernible, i.e., if it is possible to tell apart human- and machine-generated translations; ii) Whether different translation styles are automatically reproducible, i.e., if LLMs can mimic a specific translator’s style when provided with some examples; and iii) Whether style distinctions are reflected in the model’s internal representations, to motivate the interest in steering approaches for personalization.
[…]
3.1 Are Personalized Translations Discernible?
Following prior work on personalization (Wang et al., 2024c; Liu et al., 2023), we train a series of [Uncaptioned image] classifiers based on multilingual XLM Transformer encoders (Conneau et al., 2020) to distinguish between H1, H2, and MT translations. If those systems can reliably separate these three classes, it suggests the presence of reasonably distinct stylistic signals differentiating them. In particular, the ability to distinguish between H1 and H2 would denote not only the possibility to discern a human-like style from human-made and automatic translations, but also a personalized style from different human translators.
[…]
Gemma 2 2B Gemma 2 9B Llama 3.1 8B
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ZS 0.10 0.69 0.08 0.71 0.08 0.70
MS 0.24 0.69 0.31 0.73 0.32 0.73
Table 2: Classifier-based personalization accuracy ([Uncaptioned image]) and Comet-based translation quality ([Uncaptioned image]) for zero-shot (ZS) and multi-shot (MS) prompting with 20 in-context examples averaged across all translators and languages.
3.2 Can LLMs Reproduce Human Translation Styles?
To confirm whether MT personalization can be achievable, we test LLMs ability to mimic the stylistic choices of a particular translator in a multi-shot (MS) prompting setup. For each translator available across tested novels, we provide the model with 20 in-context examples selected from the original pool of translated paragraphs by that translator, asking it to generate a consistent translation. We compare MS results with the default zero-shot (ZS) prompting without any example from the translator to quantify the effect of in-context examples. Table 2 presents results for personalization accuracy, automatically evaluated using our high-scoring [Uncaptioned image] classifiers from the previous section; and translation quality, estimated via the widely used [Uncaptioned image] Comet MT metric (Rei et al., 2020). The proportion of outputs categorized as matching the translator’s style is increased two- to four-fold following MS prompting, suggesting that LLMs can employ implicit clues in small sets of user examples to produce personalized translations. Stable scores for [Uncaptioned image] Comet also confirm that translation quality is maintained during style adaptations.
3.3 Finding Personalization Information in LLM Representations
Figure 2: Probing classifier performance on the human translation detection task across Gemma 2 2B layers. Activations in intermediate layers are found to capture translation style information with high precision, with layer 13 performing best across all tested languages.
In light of these results, we set out to test where the model encodes information reflecting a stylistic shift when style-appropriate examples are provided. To this purpose, we train linear probes (Belinkov, 2022) using model activations as input features to predict the style label (MT, H1, or H2) that the style [Uncaptioned image] classifier (from Section 3.1) would assign to the eventual translation, based purely on the prompt’s internal representation. Probing accuracy is measured by testing their accuracy in predicting the classified outcome before the generation, using only the prompt representation formed by the model. Given a test set of human-translated paragraphs, we train our probes on a set of examples using an MS prompt with 20 in-context examples. The set is balanced between prompts showcasing personalization with gold in-context examples from a human translator, and non-personalized prompts with MT-generated examples previously produced by the same tested model in a ZS setup. Test examples are selected from the respective novels to ensure for the [Uncaptioned image] classifier prediction shifts from MT in the ZS setting to the style of in-context examples when MS is used, signaling a causal influence of demonstrations on output personalization. [^5] Examples are resampled for every test paragraph to prevent the probe from overfitting on spurious prompt features. This balanced setup prevents the leaking of task information, e.g., number of in-context examples, to learned probes, ensuring that stylistic differences among human and MT-generated in-context examples are the sole factor determining differences in model activations. We focus specifically on Gemma models, extracting activations after the attention block at each model layer for the last token of the prompt, which was previously shown to encode key task-relevant information (Hendel et al., 2023; Todd et al., 2024; Scalena et al., 2024)Figure 2 reports probe accuracies across all Gemma 2 2B layers, with results for the 9B model reported in Appendix C. We find a peak in probe accuracy of $\sim$95around intermediate model layers, suggesting that these layers encode stylistic information with near-perfect precision. [^6] We find probes for layers 13 and 21 to perform best for the 2B and 9B models, respectively. These results confirm that personalization is discernible from LLMs’ internal representation, motivating our experiments towards the design of inference-time interventions to steer models towards personalized MT outputs.
4 Methods
We begin by introducing our the prompting and steering methods that we use as baselines and outline our own proposed SAE-based steering approach for personalized translation.
4.1 Prompting Baselines
Zero-Shot (ZS).
The ZS setup used in our main experiment correspond to the one from Section 3.1, in which the model is simply asked to produce a translation with no conditioning from examples or explanations towards the target translation style. We use this setting to establish a baseline style and translation quality performance for the models.
Zero-Shot Explain (ZS-Expht and ZS-Exppt).
Building upon the ZS setting, we experiment with a prompting strategy where LLMs are provided with detailed explanations (Exp) of the most salient elements that characterize the desired translation style. We obtain such descriptions by prompting a capable proprietary model, GPT-4o OpenAI (2024), with 20 translations matching the desired style, asking it to synthesize a set of guidelines to matching the examples. We evaluate two contrastive variants of this approach, providing GPT-4o with either MT examples (ZS-Expht) or alternative human translations (ZS-Exppt) alongside examples matching the desired style, and asking to describe what characterizes the latter compared to the former. To avoid data leakage, all generated explanations are manually reviewed to ensure they do not contain any verbatim content or direct excerpts from the input examples. [^7] Details on the prompt templates are in Appendix B.5. Tested models are then prompted with GPT-4o explanations in a ZS setting, to verify whether interpretable directives synthesized from a set of examples matching the desired behavior can produce reliable personalization results.
Multi-Shot (MS).
Following Section 3.2’s findings, we adopt the same MS setup using 20 in-context translation examples matching the style of a target human translator (H1 or H2).
4.2 Steering Baselines
Activation Addition (ActAdd).
ActAdd is a simple yet effective technique for steering language models. We employ the standard contrastive approach by Rimsky et al. (2024); Scalena et al. (2024) to extract two sets of style-relevant ($\left\{z\right\}^{+}$) and default ($\left\{z\right\}^{-}$) activations from a given model layer using 20 in-context examples demonstrating default behavior (MT) and the desired behavior (H1 or H2 translations), respectively. We then compute the average $\Delta$ steering vector between the two sets of activations, scale it by a factor of $\alpha=2$ which was found effective by previous research (Scalena et al., 2024) and apply it additively to the same model layer during inference.
Representation Fine-tuning (ReFT).
As an alternative to traditional weight-based Parameter-Efficient Fine-Tuning (PEFT) techniques it learns task-specific interventions applied directly to model activations at inference time (Wu et al., 2024). We apply ReFT to the same personalization-relevant layers identified in Section 3.3 and limit confounding factors by tuning ReFT interventions with the set of 20 examples used for MS prompting.
4.3 Contrastive SAE Steering
Given a set of LLM activations $z_{l}\in\mathbb{R}^{d}$ sourced from the output of layer $l$, where $d$ is the model hidden size, a sparse autoencoder learns a projection to an overcomplete latent space $\mathbb{R}^{m}$, with $m\gg d$, from which a reconstructed version $z^{*}_{l}$ of the original activations is then produced with minimal information loss:
$z^{*}_{l}=\text{SAE}(z_{l})=h(z_{l})W^{l}_{\text{dec}}+b_{\text{dec}}$
[…]
where $W^{l}_{\text{enc}},W^{l}_{\text{dec}}$ are the SAE encoder and decoder modules, and $b_{\text{enc}},b_{\text{dec}}$ are bias values. Our primary interest lies in the sparse latents $h(z_{l})\in\mathbb{R}^{m}$ learned by the SAE encoder, which were empirically found to capture monosemantic and interpretable properties of model inputs.
Contrastive prompt setup
Given a set of paragraphs $\mathcal{D}$ for a novel in the par3 dataset, each instance in it is a tuple:
$\mathcal{D}=\left\{\left<s,\text{{\color[rgb]{0.2734375,0.38671875,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.38671875,0.3828125}H1}},\text{{\color[rgb]{0.51953125,0.73828125,0.61328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.51953125,0.73828125,0.61328125}H2}},\text{{\color[rgb]{0.796875,0.47265625,0.2265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.796875,0.47265625,0.2265625}MT}}_{\text{model}}\right>\right\}$
[…]
capturing respectively the personalized style of interest, and baseline properties of the task. Similarly to the ZS-Exp setup from Section 4.1, we explore two $\mathcal{D}^{-}$ configurations using either $e^{-}=\;$MT (SAE Cont.ht) or $e^{-}=\;$H2  (or H1, if H2 is the personalization target) to assess the effect of baseline choice in steering effectiveness.
Feature extraction
First, we gather activations $z^{+}_{l}$ and $z^{-}_{l}$ by prompting the model with inputs from the two contrastive sets $\mathcal{D}^{+}$ and $\mathcal{D}^{-}$. Activations are extracted at the last prompt token position from its most informative layer, as identified in Section 3.3. Activations are then converted into sparse latent representations $x^{+}=h(z^{+})$ and $x^{-}=h(z^{-})$, with $x^{+},x^{-}\in\mathbb{R}^{m}$ by the SAE encoder. This procedure is repeated across 20 contrastive examples, resulting in two collections of SAE latent vectors for positive/negative examples:
$\mathcal{X}^{+}$ =$=\left\{x^{+}_{1},x^{+}_{2},\dots,x^{+}_{20}\right\}$
$\mathcal{X}^{-}$ =$=\left\{x^{-}_{1},x^{-}_{2},\dots,x^{-}_{20}\right\}$
Relevance-based Feature Selection
To identify discriminative features for personalization in the large set of latents, we employ an information-theoretic approach adapted from Zhao et al. (2025). For each of the inputs, we identify the subset of size $n<m$ including only the SAE active features, i.e. latent dimensions for which the logit is $>0$. We consider logit values in this subset as instances of a random variable $X_{i}\in x$, and calculate the mutual information $I(X_{i},Y)$ between each feature $X_{i}$ and the target binary variable $Y=\left\{+,-\right\}$ corresponding to the style of the provided examples (personalized or non-personalized). A higher $I(X_{i},Y)$ indicates that the $i$-th feature is more informative for discriminating between personalized and default inputs, and can hence be used for steering. A representative sample of 40 latents showing the highest mutual information scores for both personalized ($\left\{X_{i}\right\}^{+}$) and non-personalized ($\left\{X_{i}\right\}^{-}$) examples is selected using this procedure.  [^8] By contrast, traditional SAE-based steering methods only employ features associated with the positive class (Chalnev et al., 2024; Arditi et al., 2024). For every selected latent, we compute its expected logit when personalization is present or absent in provided examples, i.e. $\mathbb{E}^{+}[X_{i}]$ and $\mathbb{E}^{-}[X_{i}]$.
Inference-time intervention
Finally, activations are steered by setting selected latents to their expected value whenever their observed score is below (for the promoted personalized case) or above (for the demoted non-personalized case) the pre-computed average. [^9] Algorithm 1 provides a summary of our SAE-based steering approach. Hence, in the SAE Cont.ht setting we enhance the features relevant to a target personalized style, e.g. $\left\{X_{i}\right\}^{\text{{\color[rgb]{0.2734375,0.38671875,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.38671875,0.3828125}H1}}}$ for H1, and suppress the features $\left\{X_{i}\right\}^{\text{{\color[rgb]{0.796875,0.47265625,0.2265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.796875,0.47265625,0.2265625}MT}}}$, corresponding to the model’s default MT. In SAE Cont.pt, instead, we promote the same H1-related latents while suppressing $\left\{X_{i}\right\}^{\text{{\color[rgb]{0.51953125,0.73828125,0.61328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.51953125,0.73828125,0.61328125}H2}}}$ to steer the model towards H1 personal style. Additionally, we modulate the magnitude of the resulting vector with an $\alpha$ coefficient, which was found to play an essential role in steering effectiveness in previous research (Scalena et al., 2024; Ferrando et al., 2025)
5 Experiments
5.1 Setup
Model selection
We evaluate our methods on the same three models used for our preliminary evaluation of Section 3. Our selection is guided by the availability of open-source pre-trained SAEs, which can be otherwise computationally expensive to train. For Gemma models, we employ SAEs from the GemmaScope suite (Lieberum et al., 2024); for the Llama 3.1 model we employ the SAE released by (McGrath et al., 2024). GemmaScope SAEs are available for every model layer, enabling us to steer Gemma models on their most informative layers for the task, which we identified in Section 3.3. On the contrary, a single SAE for the 19th layer is available for Llama, hence limiting our evaluation of SAE steering and potentially producing sub-optimal steering result for that model.
Gemma 2 2B Gemma 2 9B LLaMA 3.1 8B
H P ${}_{\textsc{flip}}$ [Uncaptioned image] H P ${}_{\textsc{flip}}$ [Uncaptioned image] H P ${}_{\textsc{flip}}$ [Uncaptioned image]
ZS 0.21 0.10 0.05 0.69 0.15 0.08 0.04 0.71 0.24 0.08 0.05 0.70
ZS-Expht 0.30 0.22 0.16 0.68 0.41 0.22 0.18 0.72 0.56 0.23 0.21 0.69
ZS-Exppt 0.20 0.14 0.69 0.23 0.19 0.73 0.30 0.26 0.70
MS 0.37 0.24 0.16 0.69 0.48 0.31 0.27 0.73 0.58 0.32 0.28 0.73
ActAdd 0.27 0.22 0.12 0.67 0.32 0.24 0.20 0.70 0.55 0.36 0.28 0.70
ReFT 0.31 0.22 0.18 0.70 0.46 0.34 0.27 0.67 0.53 0.38 0.26 0.70
SAE Cont.ht 0.39 0.27 0.19 0.70 0.46 0.33 0.29 0.72 0.59 0.31 0.27 0.72
SAE Cont.pt 0.27 0.18 0.69 0.35 0.29 0.73 0.33 0.28 0.72
[…]
Metrics
[…]
To ensure that our interventions do not result in a degradation of overall translation quality, we also employ [Uncaptioned image] Comet [^10] Unbabel/wmt22-comet-da (Rei et al., 2020) using the personalized translation as reference.
The $\alpha$ trade-off
We begin by verifying the optimal steering intensity $\alpha$ for our SAE steering technique. We primarily focus on results from Gemma 2 2B, for which we ran a comprehensive sweep over all relevant hyperparameters. [^11] Larger models were evaluated using a subset of the best-performing configurations. Details in Appendix C. Figure 3 illustrates the influence of $\alpha$ on MT personalization accuracy and fluency averaged across all translators for all tested languages. For values of $\alpha\leq 3$, performance remains close to that of the MS baseline, indicating that the contrastive method is effectively isolating latents associated with human-like style. As $\alpha$ increases, performance generally exceeds the MS approach, achieving greater control and flexibility in guiding the model’s output with next to no impact of translation quality. However, for $\alpha\geq 10$, we observe a major degradation in [Uncaptioned image] Comet, suggesting an important drop in translation fluency.
Following Ferrando et al. (2025), which also employ SAEs for steering, we experiment with very high alpha values (up to 150), finding the [Uncaptioned image] classifier’s H accuracy approaching 100% for some languages. While this indicates that the contrastive steering is aggressively optimizing toward classifier preferences (Figure 4), the consequent drop in [Uncaptioned image] Comet scores reveals a steep decline in translation quality, often resulting in incoherent or nonsensical generations from a human perspective. [^12] A qualitative evaluation is provided in Appendix D. Ultimately, we identify $\alpha=5$ as an appropriate steering intensity to balance personalization and fluency, and employ it for our main evaluation.
Figure 3: Personalization P and Refer to caption Comet across various steering intensity $\alpha$ for SAE Cont.ht on Gemma 2 2B. The performance of prompting baselines (ZS, MS, Exp) is also reported. Results show a trade-off between steering intensity and translation quality.
Figure 4: Refer to caption Comet and H accuracy across $\alpha$ steering intensity values for Gemma 2 2B, showing a major drop in translation quality for very high intensities ($\alpha\geq 50$).
5.2 Results and Discussion
Table 3 presents performances of tested models across prompting and steering setups, averaged across all languages and personalization targets (H1 and H2 for each language). We find that our SAE Cont.ht and SAE Cont.pt methods generally achieve the best trade-off between personalization accuracy and translation quality, especially for the smaller Gemma 2 2B model. This could be due to the larger models’ superior ability to incorporate in-context information naturally, reducing the relative benefit of explicit steering.
Which contrastive setup is better?
Comparing the two contrastive setups (ht and pt) for the ZS-Exp and SAE Cont. methods, we find that using different human demonstrations as a contrastive baseline in pt generally produces better results for larger models. As for general performance, we conjecture this could be due to the larger models’ improved ability to disentangle personalization-critical factors without explicit guidance. For the smaller Gemma 2 2B, the difference between the two approaches is minimal, suggesting the model cannot fully exploit the examples’ differences.
Do SAE Steering and MS Prompting Impact Activations in a Similar Way?
Since SAE-based approaches perform on par or better than MS, we set out to investigate whether the two methods result in a similar impact on model representations. We collect the modified activations $z_{\text{steer}}$ obtained from the SAE Cont.ht steering setting and evaluate them using the probing classifier trained on MS-conditioned activations we introduced in Section 3.3 for detecting personalization information. Table 4 shows probe accuracy in detecting the positive impact of SAE steering across the three possible outcomes of the steering procedure. We find that the probe corresponding to the SAE layer effectively distinguishes between activations corresponding to successful and unsuccessful SAE steering, despite having been exposed only to MS conditioning during training. This includes both instances for which the classifier prediction is flipped after steering (MT $\rightarrow$ H*), and settings where the conditioning fails (MT $\rightarrow$ MT). In settings where the original output already matches human style (H* $\rightarrow$ H*), the probe obtains lower accuracy with broader confidence intervals, denoting higher uncertainty. These findings suggest that the SAE’s latents we extract through our contrastive method are meaningfully connected to the stylistic patterns embedded in the multi-shot examples, providing evidence that our intervention influences the internal representations of the model, aligning them to the natural effect of the MS approach.
MT $\rightarrow$ H* MT $\rightarrow$ MT H* $\rightarrow$ H*
Gemma 2 2B 0.94
±$\pm 0.01$
0.07
±$\pm 0.02$
0.72
±$\pm 0.15$
Gemma 2 9B 0.93
±$\pm 0.02$
0.12
±$\pm 0.10$
0.68
±$\pm 0.19$
Table 4: Probing accuracy on SAE Cont.ht-steered activations averaged across languages and H1/H2 translators. Probes trained on MS activations reliably detect the impact of SAE Cont. steering over model generations, suggesting similar mechanisms.
6 Conclusion and future work
We conducted a broad evaluation of various prompting and steering approaches for personalizing LLM-generated translations. Our evaluation targets a practical, real-world application of literary translation and addresses the underexplored challenge of steering LLM generations in a linguistically rich and stylistically sensitive domain. Through comprehensive evaluation across multiple languages, novels, and models, we demonstrate that our proposed SAE-based approach outperforms prompting and alternative steering techniques.
Although faithfully replicating individual human translation styles remains a highly challenging task, our approach achieves strong alignment with human translation quality, as reflected in both general human-likeness and translator-specific personalization metrics. These results highlight the method’s robustness and its potential to support high-fidelity translation workflows in real-world settings. Concretely, these results have important implications in the development of personalized MT systems based on LLMs. In particular, the notable effectiveness of our proposed approach on smaller models might enable MT personalization in lower-resource settings, easing further research on how personalization information is encoded and produced by language models.
Future work will focus on improving the interpretability of the learned SAE latents, to uncover their learned monosemantic features. Additionally, we plan to explore the integration of our method with larger language models, where increased capacity may further enhance the precision and fluency of personalized translations.
Limitations
While our work demonstrates the potential of steering LLMs for MT personalization using sparse autoencoders, we acknowledge several limitations.
Firstly, our findings’ generalizability is constrained by the scope of our experiments. We focused on literary translation into English from seven specific source languages and evaluated three LLMs of relatively small size. Consequently, the observed effectiveness of SAE-based steering and the identified optimal layers for intervention may not directly transfer to other language pairs, significantly different model architectures or sizes, or distinct domains beyond literary texts. Further research is needed to assess the robustness of our approach across a broader range of linguistic and modeling contexts.
Secondly, the computational overhead associated with sparse autoencoders presents a practical challenge. Although we utilized pre-trained SAEs in our study, the initial training of these components is resource-intensive. This could limit the accessibility and scalability of our proposed method, particularly for researchers or practitioners with limited computational resources or when frequent retraining for new models or tasks is required. The current availability of pre-trained SAEs also restricts model choice, as seen with the Llama 3.1 8B model where an SAE was only available for a potentially sub-optimal layer.
Finally, our investigation primarily focused on downstream performance and the impact of various personalization strategies on model representations. However, we did not pursue a mechanistic understanding of the "personalization circuits" within the LLMs. Future work could adopt a more fine-grained, mechanistic interpretability approach to study how specific SAE latents or combinations thereof encode and manipulate nuanced stylistic features, thereby providing deeper insights into the underlying processes of LLM personalization.
Appendix A List of novels used
To ensure a diverse and representative evaluation, we select novels spanning a variety of linguistic families and cultural backgrounds. Our dataset includes Romance languages such as Italian (Pinocchio) and French (Around the World in Eighty Days), as well as Germanic languages like Dutch (The Diary of a Young Girl) and German (Beware of Pity). To evaluate our setup on non-Latin scripts and distinct linguistic structures, we also include Russian (Crime and Punishment), Japanese (No Longer Human), and Chinese (Dream of the Red Chamber). Table 5 summarizes the number of paragraphs employed in the evaluation of each language.
Lang Novel name Train Val Test ICL
IT Pinocchio 745 82 107 20
FR Around the World in Eighty Days 829 92 120 20
NL The Diary of a Young Girl 769 85 110 20
DE Beware of Pity 606 67 96 20
RU Crime and Punishment 1517 168 224 20
JA No Longer Human 652 40 81 20
ZH Dream of the Red Chamber 694 76 92 20
Table 5: Each number corresponds to a single instance in the dataset. When the dataset is used for training, each instance is associated with three distinct labels - H1, H2, and MT - thus the total number must be multiplied by three.
Appendix B Experiments reproducibility
In this section, we provide every parameter we use for the reproducibility of our experiments setups.
B.1 Base prompt
We use the same prompt template across all methods: ZS (which corresponds to the original model translation), ZS-Expht, ZS-Exppt (detailed in Section B.5), MS, ActAdd, ReFT, and SAE-based contrastive setups. This prompt, shown in Listing LABEL:lst:explain-final, instructs the model to translate the source sentence while explicitly preventing it from adding any explanations about the translation process. Since all test models are Instruction Tuned, we utilize their native chat templates to preprocess the input accordingly. For multi-shot examples, the user and assistant turns are repeated for each example, always using the same prompt structure.
B.2 Classifier training
All classifiers are fine-tuned from the xlm-roberta-large model [^13] FacebookAI/xlm-roberta-large, using a linear classification head. Training is conducted for 6 epochs with a learning rate of 2e-5 and a batch size of 32, selecting the best model checkpoint based on validation accuracy.
Training data only includes generations from models and the translator without any source text. It is also perfectly balanced, as each paragraph provides one instance for all three labels: H1, H2, and MT. The total size of the training set varies depending on the number of paragraphs in the chosen novel. On average, we obtain approximately 830 instances, resulting in a total of around 2,490 labeled examples for training (see 5). Validation and test sets are strictly held out and never seen during training. Additionally, they do not include the small 20-example subsets used in the MS, ZS-Expht, ZS-Exppt, SAE Cont.ht, and SAE Cont.pt setups.
B.3 ReFT training
ReFT training was conducted using the PyReFT toolkit from the original authors141414stanfordnlp/pyreft. We applied the intervention at the same hook point used by other steering methods - specifically, the layer output corresponding to the residual stream at the selected layer. The training configuration includes a low rank dimension of 4, lora alpha set to 32, and a lora dropout of 0.05. ReFT was trained on the same 20 prompts used in the MS setup, for a total of 100 epochs.
B.4 SAE Cont.ht and SAE Cont.pt
We use the NNsight library (Fiotto-Kaufman et al., 2024) to extract and manipulate model activations for all steering experiments. The source code is publicly available in the repository linked in the main body of this paper. For consistency, we use the same set of contrastive examples employed in the MS approach.
Algorithm 1 outlines the procedure for latent-based steering. It enhances features identified as relevant to personalization while simultaneously suppressing those negatively correlated with the task.
Input: Input activation $z$, SAE model sae, target latents expected value $\mathbb{E}^{+}[X_{i}]$, contrast latents expected value $\mathbb{E}^{-}[X_{i}]$, steering coefficient $\alpha$
Output: Steered activation $z_{\text{new}}$
$x\leftarrow\texttt{sae.encode}(z)$;
$m\leftarrow\text{length}(x)$;
for $i\leftarrow 1$ to $m$ do
       if $\mathbb{E}^{+}[X_{i}]>x[i]$ then
             $x[i]\leftarrow\mathbb{E}^{+}[X_{i}]$
      if $\mathbb{E}^{-}[X_{i}]<x[i]$ then
             $x[i]\leftarrow\mathbb{E}^{-}[X_{i}]$
$z_{\text{steer}}\leftarrow\alpha\cdot\texttt{sae.decode}(x)$;
return $z_{\text{steer}}$
Algorithm 1 Contrastive SAE Steering
B.5 ZS-Expht and ZS-Exppt
For both the ZS-Expht and ZS-Exppt setups, we used GPT-4o (June 2025) to generate explanations detailing the stylistic differences between a base translation and a target human translation. The prompt template used for this task is shown in Listing LABEL:lst:explain-template, using the same 20 examples as in the MS, SAE Cont.ht, and SAE Cont.pt setups.
All outputs were manually inspected to ensure no verbatim excerpts from the provided examples were present, avoiding any risk of data leakage. Example outputs for different novels are shown in Listing LABEL:lst:explain-example.
Finally, these generated guidelines are used to prompt the evaluated models, following the template shown in Listing LABEL:lst:explain-final.
Listing 1: Prompt template used to get GPT 4o explanation using translation examples.
Objective: Identify stylistic choices in translations for personalization purposes.
You will be provided with a source text, a standard translation, and a target translation by a specific translator whose style we want to emulate.
Your task is to analyze the Target translation by comparing it to the Base translation and the Source text’.
Identify and list the distinctive stylistic patterns, choices, and preferences exhibited in the Target translation.
These stylistic cues should help another translator (or an AI) to adapt their translations to match the style of the target translator.
Source text: <source text here>
Base translation: <MT / H- text here>
Target translation: <H+ translation here>
<… repeat Source, MT and Target>
Please extract a concise list of key stylistic cues. Focus on aspects such as vocabulary choices, sentence structure, tone and register, handling of cultural nuances, punctuation/formatting preferences and overall creativity.
Output a short list of stylistic cues as bullet points. Write the list as if you were directly giving the guidelines to the translator and avoid using specific examples.
Listing 2: Examples of explanation obtained from GPT 4o when comparing different translations from different novels.
When comparing H1 and MT for Beware of Pity (German):
- Maintain a tone that is professional, thoughtful, and subtly persuasive.
- Avoid overly technical jargon unless necessary; explain specialized terms briefly if used.
- Preserve the authors voice, keeping a balance between academic rigor and narrative engagement.
- Ensure smooth transitions between sentences and paragraphs to support coherent argumentation.
- Translate idiomatic expressions in a way that retains their intended effect, even if the wording differs.
When comparing H1 and H2 for Beware of Pity (German):
- Keep the tone warm, welcoming, and direct - avoid overly formal or distant language.
- Use active voice wherever possible to maintain energy and engagement.
- Avoid idioms or expressions that may not translate culturally; aim for universal accessibility.
- Maintain consistent tone and register throughout, adapting to the intended audiences familiarity with the subject.
- Respect the rhythm and structure of the original, but feel free to adjust sentence length for readability.
When comparing H1 and H2 for Crime and Punishment (Russian):
- Prefer dynamic over formal vocabulary: Opt for vivid or emotionally charged words when available.
- Add tonal nuance and emotional shading: Enrich dialogues and narration with subtle shifts in tone, especially sarcasm, understatement, or irony, to match character voice or mood.
- Use contractions and familiar phrasing: Employ contractions and relaxed expressions to preserve spoken character.
- Expand or rephrase for clarity and voice: Dont hesitate to slightly reword or elaborate if it strengthens tone, clarifies intent, or enhances character differentiation.
- Favor rhythmic, flowing sentence structure: Break long, formal sentences into multiple shorter clauses or use punctuation (dashes, ellipses) for dramatic or emotional effect.
- Reflect subtle character dynamics: Infuse lines with interpersonal undertones (like defiance, deference, or sarcasm) that may not be explicit in the original.
- Preserve or recreate emotional tension: Use word choice and pacing to sustain psychological nuance, unease, or irony.
- Use expressive punctuation and formatting: Favor dashes, ellipses, and italic like emphasis (through word placement) to reflect emotional cadence or interruptions.
Listing 3: Zero shot template template when prompting language models with different setups
Translate the following sentence between the angular parentheses into English.
if setup == ZS-Exp@
Follow the following guidelines when translating:
<explanations here>
The original sentence is: <source text>.
Remember to write only the translation, without any additional text or explanation.
Appendix C All models results
C.1 Classifiers
We show in Table 6 results for every [Uncaptioned image] classifier trained for each model and for each language.
Figure 5: Probing classifier performance on the human translation detection task accross Gemma 2 9B layers. For our experiments we select layer 21 as the optimal intervention point for our steering approaches.
Lang. Gemma 2 2B Gemma 2 9B Llama 3.1 8B
DE 0.89 0.90 0.84
RU 0.92 0.90 0.91
ZH 0.99 0.98 0.98
IT 0.78 0.85 0.80
NL 0.79 0.78 0.82
FR 0.88 0.87 0.90
JA 0.76 0.79 0.76
Table 6: Accuracy of model- and language-specific 3-way (MT, H1, H2) classifiers on balanced held-out sets for every language. Random baseline: $0.33$.
C.2 Probing experiments
We report in Figure 5 the probing classifier performance for Gemma 2 9B. We select the 21th layer as the most informative one to make our intervention. Results for the smaller model Gemma 2 2B are shown in the main body of the paper in Figure 2.
C.3 Prompting and steering results
We present detailed plots of the results for each novel across the three evaluated models in Figure 6 (Gemma 2 2B), Figure 7 (Gemma 2 9B), and Figure 8 (Llama 3.1 8B). These plots display the performance of all evaluated methods, reporting the three submetrics: H accuracy (general human-likeness), P accuracy (translator-specific accuracy), and ${}_{\textsc{flip}}$ (personalized flip accuracy), alongside the corresponding [Uncaptioned image] Comet scores measuring translation quality.
Figure 6: Results for every language on Gemma 2 2B.
Figure 7: Results for every language on Gemma 2 9B.
Figure 8: Results for every language on Llama 3.1 8B.
Appendix D Examples from dataset and different approaches
We present in Tables 7 and 8 a selection of examples from two different languages, showcasing outputs from each of the tested setups. For each example, we also report the corresponding classification label predicted by the [Uncaptioned image] classifier and the associated [Uncaptioned image] Comet score. Additionly table 9 shows some examples of models generating output aligned with the Human translator according to the [Uncaptioned image] classifier but with a low [Uncaptioned image] Comet score corresponding to an almost unreadable output due to extreme $\alpha$ values.
Source (ZH) 遂又往後看時,只見畫著一張弓,弓上掛著香櫞。也有一首歌詞云:「二十年來辨是非,榴花開處照宮闈。三春爭及初春景,虎兕相逢大夢歸。」 [Uncaptioned image] classifier [Uncaptioned image] Comet
H1 Forthwith, therefore, he prosecuted a further perusal of what came next, when he caught sight of a picture of a bow. On this bow hung a citron. There was also this ode: Full twenty years right and wrong to expound will be thy fate! What place pomegranate blossoms come in bloom will face the Palace Gate! The third portion of spring, of the first spring in beauty short will fall! When tiger meets with hare thou wilt return to sleep perennial. H1 -
H2 The picture that followed was of a bow with a citron hanging from it, followed by what looked like the words of a song: You shall, when twenty years in life’s hard school are done, In pomegranate-time to palace halls ascend. Though three springs never could with your first spring compare, When hare meets tiger your great dream shall end. H2 -
Gemma 2 2B When he turned back to look, he saw a bow hanging on it, with a plum blossom. There was also a poem that said: "For twenty years, I have been trying to distinguish right from wrong, and the plum blossoms bloom in the palace. In the spring, the tigers and the rhinoceroses meet, and they dream of returning home." MT 0.49
Gemma 2 9B Looking back, he saw a painting of a bow with a quiver hanging on it. There was also a line of lyrics that said: F̈or twenty years, I have distinguished right from wrong, and the blossoming of the pomegranate flowers illuminates the imperial court. Three springs have passed, and now it is the beginning of spring again, like a dream where tigers and rhinoceroses meet. MT 0.56
Llama 3.1 8B As I looked back, I saw a painting of a bow with a censer on it. There was also a poem that said: T̈wenty years of distinguishing right from wrong, the lotus flower blooms in the palace. Three springs competing with the scenery of the first spring, the tiger and the rhinoceros meet and return to their dream." MT 0.53
ZS-Expht And as he turned his gaze back upon the scene once more, there was displayed an archer’s bow strung tautly; upon its surface hung a fragrant plum blossom. A poem also recited thus spake: "For twenty years I have sought truth amidst strife, where blooming durian flowers illuminate the palace walls. In spring’s embrace, they vie for beauty, tigers clashing in dreams of return." MT 0.60
ZS-Exppt And as he turned his gaze back upon the scene once more, there was displayed an archer’s bow strung tautly; upon its surface hung a fragrant plum blossom. A poem also recited thus: "For twenty years I have sought truth amidst strife, where blooming durian flowers illuminate the palace walls. In spring’s embrace, they vie for beauty, tigers clashing in dreams of return." MT 0.61
MS Then, turning around, he noticed a picture depicting a bow hanging on its string, adorned with incense sticks. A verse read: "For twenty years I’ve tried to discern truth from falsehood, and wherever fragrant plum blossoms bloom, the palace gates shine brightly." H1 0.65
ActAdd When he turned back to look, he saw a bow hanging on it, with a plum blossom. There was also a poem that said: "For twenty years, I have been trying to tell right from wrong, and the plum blossoms bloom in the palace. In spring, the tigers and rhinoceroses meet, and they dream of going back home." MT 0.53
ReFT When he turned back to look, he saw a bow hanging there, adorned with a plum blossom. There was also a poem that said: "For twenty years, I have tried to distinguish right from wrong, and plum blossoms bloom within the palace. In spring, tiger and rhinoceros meet, dreaming of their return home." MT 0.58
SAE Cont.ht
   $\alpha=1$ When he turned back to look, he saw a bow hung upon it, with a plum blossom. There was also a verse which read: "For twenty years, I have struggled to discern truth from falsehood, and the plums blossomed where the imperial court resided. During the three springs, the tiger and rhinoceros met, dreaming of their return." MT 0.62
   $\alpha=5$ He looked back and there hung a bow adorned with a plum blossom. A verse read: “Twenty years I’ve struggled for truth, where fragrant plums bloom near the imperial court. Spring arrives, tiger and rhino clash, dreaming of their return. MT 0.62
   $\alpha=150$ The when the veryShellTruth lookWarmTheTheTheTigeratPoint […] MT 0.11
SAE Cont.pt
   $\alpha=5$ He looked back and saw a bow adorned with fragrant citron. A verse declared: "Twenty years I’ve weighed right and wrong; where pomegranate blossoms bloom, they light the palace gates. Spring’s full bloom cannot match early spring’s grace. When tiger meets rhino, one returns to the great dream." H1 0.65
Table 7: Example paragraphs from “Dream of the Red Chamber” (ZH $\rightarrow$ EN) translated with H1personalization. Setup outputs are from Gemma 2 2B; only MS and SAE Cont.pt successfully flip the [Uncaptioned image] classifier’ prediction.
Source (IT) Intanto la stalla fu chiusa e Pinocchio rimase solo: e perché erano molte ore che non aveva mangiato, cominció a sbadigliare dal grande appetito. E, sbadigliando, spalancava una bocca che pareva un forno. [Uncaptioned image] classifier [Uncaptioned image] Comet
H1 When the stable door was closed, Pinocchio was left alone. And since he hadn’t eaten in many hours, he began to yawn from hunger. And when he yawned, his mouth opened as wide as an oven. H1 -
H2 The stable was then shut and Pinocchio was left alone. He had not eaten for many hours, and he began to yawn from hunger. And when he yawned he opened a mouth that seemed as wide as an oven. H2 -
Gemma 2 2B At that moment the stable was closed and Pinocchio was left alone, and because it had been many hours since he had eaten, he began to yawn from a great hunger. And, yawning, he opened a mouth that seemed like a furnace. MT 0.79
Gemma 2 9B Meanwhile, the stable was closed and Pinocchio was left alone: and because it had been many hours since he had eaten, he began to yawn from his great hunger. And, yawning, he opened a mouth that looked like an oven. MT 0.80
Llama 3.1 8B Meanwhile the stable was closed and Pinocchio was left alone: and because it was many hours since he hadn’t eaten, he began to yawn from great hunger. And, yawning, he opened a mouth that seemed like an oven. MT 0.80
ZS-Expht Meanwhile, the stable door was shut, leaving Pinocchio alone. His tummy rumbled because it had been hours since his last meal, and he started yawning widely from his hunger. With each yawn, his mouth opened like an oven. MT 0.78
ZS-Exppt Meanwhile, the stable door was shut, leaving Pinocchio all alone. His tummy rumbled because it had been hours since his last meal, and he started yawning widely from sheer hunger. With each yawn, his mouth gaped open like an oven. H2 0.80
MS Meanwhile, the stable door was shut and Pinocchio was left alone. Since it had been hours since he last ate, he began yawning from hunger. Each yawn opened his mouth wide enough to look like an oven. H2 0.81
ActAdd Meanwhile, the stable was closed and Pinocchio remained alone. Because many hours passed without eating, he started yawning from very big hunger. And while yawning, he opened his mouth wide like an oven. MT 0.75
ReFT Meanwhile, the stable was closed and Pinocchio stayed alone. Because many hours had passed since he last ate, he began to yawn from big hunger. And, yawning, he opened a mouth that looked like an oven. MT 0.77
SAE Cont.ht
   $\alpha=1$ Meanwhile, the stable door was shut and Pinocchio was left alone. Because it had been many hours since he had eaten, he started to yawn with intense hunger. And, as he yawned, his mouth gaped open like an oven. MT 0.79
   $\alpha=5$ Meanwhile, the stable door shut, leaving Pinocchio all alone. Because it had been many hours since he’d eaten, he started to yawn with intense hunger. As he yawned, his mouth gaped open, looking like an oven. H2 0.82
   $\alpha=50$ He He "It The … You It "[They They) THE…I I HE IT [There There THEYRE WE "A A FOR "[…] MT 0.14
SAE Cont.pt
   $\alpha=5$ The stable was then shut, leaving Pinocchio all alone. He had not eaten for many hours and began to yawn from great hunger. As he yawned, he opened his mouth wide, like an oven. H2 0.84
Table 8: Example paragraphs from “Pinocchio” (IT $\rightarrow$ EN) translated with H2personalization. Setup outputs are from Gemma 2 9B; ZS-Exppt, MS, SAE Cont.ht and SAE Cont.pt (both $\alpha=5$) can flip the [Uncaptioned image] classifier prediction.
Source (FR) Cette somme vous sera restituée á votre sortie de prison, dit le juge. En attendant, vous êtes libres sous caution. [Uncaptioned image] classifier [Uncaptioned image] Comet
H1 This sum of money will be returned to you when you leave prison,’ said the judge. ’In the meantime you are free on bail. - -
ZS This amount will be returned to you upon your release from prison, the judge said. Meanwhile, you are free on bail. MT 0.79
SAE Cont.ht
   $\alpha=$ 5 This sum will be repaid to you at your departure from prison, the judge declared. In the meantime, you are released under bond. MT 0.85
   $\alpha=$ 50 #echo all ought to goodness too was put moreover goodness behaving goodness goodness goodness goodness […] H1 0.20
   $\alpha=$ 150 GURL callers goodness_Noneummy amidála Reportedaps Beaverishlor Decompiled_unset queer headquarters Headquarters twe … pocket-or callers dt behaving indoorsgetRepository callers_lazy uninsured tipostiFRING callers callers callers callers Guinness_None … foss bedroomsitud facet goodness goodness goodness […] H1 0.23
Source (DE) Nein – was Sie lieber haben, Herr Leutnant! Nur keine Zeremonien, es ist doch ganz einerlei. [Uncaptioned image] classifier [Uncaptioned image] Comet
H2 "No, no-whatever you would rather have, Lieutenant Hofmiller! Please don’t stand on ceremony, it makes no difference to us." - -
ZS No, whatever you prefer, Lieutenant! Just no ceremonies, it doesn’t matter. MT 0.76
SAE Cont.ht
   $\alpha=$ 5 No, anything at all you want, sir! Just don’t make a fuss about it, it really doesn’t matter. H2 0.79
   $\alpha=$ 50 ">I Don’t worry about that… I don’t want a ceremony for this one. It’s not important… H2 0.46
   $\alpha=$ 150 IWhenInWhatItDonIf Sometimes AIs Celebrating cerimonies… Sosir please don’t have parties ey’ […] H2 0.24
Table 9: Examples from different languages being classified as Human when using extreme $\alpha$ values.